# Multiple Instance Verification

### v1

@@ Abstract
We explore multiple-instance verification, a problem setting where a query instance is verified against a bag of target instances with heterogeneous, unknown relevancy. We show that naive adaptations of attention-based multiple instance learning (MIL) methods and standard verification methods like Siamese neural networks are unsuitable for this setting: directly combining state-of-the-art (SOTA) MIL methods and Siamese networks is shown to be no better, and sometimes significantly worse, than a simple baseline model. Postulating that this may be caused by the failure of the representation of the target bag to incorporate the query instance, we introduce a new pooling approach named “cross-attention pooling” (CAP). Under the CAP framework, we propose two novel attention functions to address the challenge of distinguishing between highly similar instances in a target bag. Through empirical studies on three different verification tasks, we demonstrate that CAP outperforms adaptations of SOTA MIL methods and the baseline by substantial margins, in terms of both classification accuracy and quality of the explanations provided for the classifications. Ablation studies confirm the superior ability of the new attention functions to identify key instances.
Keywords Multiple Instance Learning, Verification, Siamese Neural Networks, Attention, Explainability
@@ 1 Introduction
In multiple instance (MI) verification, each exemplar for machine learning consists of a pair of two objects—a query instance and a bag of target instances—and a binary class label that indicates whether or not the bag contains an instance pertaining to the same (unobservable) class as the query instance. [^1] For brevity, we refer to these components as “query”, “target bag”, “exemplar label”, and “query class”, respectively, in what follows. Figure 1(a) shows an example: a synthetic task based on an extended version of the MNIST data consisting of digits and their writers’ ID [38]. In this task, each exemplar contains multiple instances of the same digit written by different writers. The question is whether the target bag contains handwritten digits from the writer who wrote the query—and if there are any, which ones are “key instances". Other practical applications are illustrated in Figures 1(b) (signature verification) and 1(c) (fact verification) respectively. Such tasks are of practical importance because they represent settings that require “verification with noise”. While basic verification tasks involve comparing a query instance to a single target instance, real-world targets available for verification are often “noisy” in the sense that they consist of multiple candidates, some irrelevant, to be compared. Importantly, which individual target instances are relevant is unknown, possibly because obtaining labels of relevancy is too costly. Hence, verification needs to be conducted using the noisy target objects. Note that this generally involves establishing the candidate instances’ relevance to the outcome of the verification process as well—a potentially very useful byproduct.
@@
Conceptually, the MI verification problem may be characterized as a combination of two sub-problems: verification [2] and multiple instance learning (MIL) [10]. Verification is frequently formulated as a binary classification task based on a pair of inputs, predicting whether or not the paired inputs are of the same class. The standard solution is to apply Siamese neural networks [2, 6, 20], but they have no built-in mechanism to aggregate or select from multiple elements within an input, which renders them unsuitable for MI verification. On the other hand, algorithms for MIL, where bags of instances are to be classified, address the ambiguity in how a bag-level class label is related to the unknown labels of the instances in a bag [44, 3]. However, existing MIL methods do not provide a dedicated mechanism to take the information provided by a query instance into account. In this paper, we argue that new algorithms and models are required for this setting and provide theoretical and empirical evidence showing the benefit of explicitly incorporating both components—the target bag and the query instance—and modelling their relationship during learning and inference.
@@
We also stress that “multiple instance verification” and similar terms as used in domains such as biometrics [23, 32] refer to problems that differ from the setting considered in this paper. In those problems, all instances of a bag, by definition, have homogeneous unobservable labels, implying no noise or irrelevant elements in the bag for verification. This is in contrast to the verification problem with heterogeneous instance labels studied in this paper. Crucially, just as in standard MIL, an exemplar label only provides information about the existence of key instances in a bag but does not identify them. Moreover, as the query instance may belong to a class that has not been observed when the verification model was trained, it enables classification into previously unknown classes. Following the rationale of [20], this is analogous to an MI version of zero-shot classification.
@@
A straightforward approach to MI verification is to combine existing verification and MIL methods as building blocks, e.g., by creating a bag-level representation for the target bag using a state-of-the-art (SOTA) MI pooling method, and then treating this bag-level representation together with the query instance as the two inputs to a Siamese network for verification. Suitable SOTA MI pooling methods include approaches based on gated attention (“gated-attention-based MIL” [17]), the transformer decoder (“Pooling by Multi-head Attention” [21]), and the transformer encoder (e.g., “TransMIL” [30]). Alternatively, one may apply non-attention-based methods such as MI-Net [36] and bi-direction LSTM (Bi-LSTM) [35]. However, we show that MI verification based on this straightforward approach, regardless of which of the above MIL algorithms is applied, yields unsatisfactory classification performance. More specifically, a simplistic “baseline" model is shown to outperform the resulting instantiations—henceforth referred to as “benchmarks”—of the straightforward approach to MI verification.
@@
We postulate that the benchmarks perform unsatisfactorily because they do not incorporate the query instance into the bag-level representation of the target bag: the “query” concept does not exist in the underlying MIL methods. Indeed, we show theoretically, using a probabilistic interpretation of attention scores, that failure to incorporate the query instance induces less effective attention scores. Hence, inspired by the “query-key-value” attention mechanism of the transformer decoder [34], we propose a new pooling approach called “cross attention pooling” (CAP) that creates the bag-level representation of a target bag through a cross-reference attending to the query. This means the bag’s representation is dynamic because it varies when a different query instance is present, even if the bag itself does not change. In contrast, standard MIL methods create a static bag-level representation. CAP also enables the introduction of two novel attention functions, namely “distance-based attention” and “variance-excited multiplicative attention” respectively, to address the challenge that the instances in a bag may be difficult to distinguish. Both functions accommodate multi-head formulations popular in transformers.
@@
In addition to classification accuracy, the ability to identify the key instances in a target bag is also highly desirable for MI verification. This ability is commonly known as “explainability” or “interpretability” in the literature on MIL [17, 30]. Following recent work on this topic [18, 5], we quantitatively evaluate explanation quality for the methods in our study. This is possible for those datasets where ground-truth instance labels are available.
@@
Our experimental results, obtained from three different tasks and considering both classification and explanation performance, show that CAP is an effective solution to the MI verification problem. In terms of (bag-label) classification performance, it significantly outperforms the benchmarks and the baseline by large margins. Furthermore, using quantitative measures of explanation quality, we show that it identifies key instances far more accurately than both the benchmarks and the baseline. Higher classification accuracy and explanation quality are observed with different numbers of training exemplars and with varying bag sizes in the training data. We also present an ablation study showing that the components of the novel attention mechanisms responsible for delivering superior explanation quality—and thus leading to better bag-level representations—are the source of better classification performance.
@@
Our key contributions are as follows. First, we formally state the MI verification problem, and create a new solution, CAP, that outperforms SOTA MIL methods combined with Siamese networks and a simple baseline method on three tasks from the computer vision (CV) and natural language processing (NLP) domains. Furthermore, we propose two new attention functions for CAP: (a) distance-based attention, which differs from the mainstream functional forms of attention because it is neither additive nor multiplicative, and (b) variance-excited multiplicative attention, which is a new form of multiplicative attention. We show that they are better able to identify key instances in MI verification, resulting in better bag-level representations. While attention-based methods are not uncommon in MIL-related domains, research studying different types of attention mechanism and analyzing their efficacy under specific conditions is rare. Within this context, we provide insights that highlight the importance of quantitative evaluation of the explanation quality to identify spurious model performance. We find that MI models may identify key instances incorrectly but still classify the corresponding bag correctly, and vice versa. The work also contributes to the specific application areas we consider. For example, in the natural language inference (NLI) subtask of “fact extraction and verification”, the existing approach is a two-step pipeline: evidence retrieval and support verification. Our approach combines them into a single step through CAP, and thus enables the end-to-end training of the whole pipeline.
@@ 2 Related work
The most closely related existing work is research on verification using Siamese neural networks and the deep MIL literature.
@@
The verification problem, and its solution using Siamese neural networks, was first introduced by the pioneering work of [2]. [6] further explored Siamese networks for similarity metric learning. Since then, the problem settings of verification have been generalized, in both CV (e.g., [41]) and NLP (e.g., [7]) domains where Siamese networks can be applied. There are also several areas in machine learning that apply Siamese networks to solve problems similar to verification, including one-shot learning [20, 15], self-supervised and contrastive learning [4, 42], embedding pre-training [7, 28], to name a few examples.
@@
We note that there exist approaches other than Siamese networks that may also be applicable to MI verification in specific application scenarios, such as using concatenation of the query instance and the target bag in the NLP task considered in [29]. We do not consider this approach here because it needs to be used jointly with other NLP techniques and cannot be trained end-to-end. Within the scope of this paper, we only consider Siamese networks, which provide a more generic and domain-agnostic approach.
@@
As for MIL, starting from the seminal work of [10], it forms an important area called “weakly supervised learning” (see [44]). Loosely speaking, depending on whether a method (1) explicitly creates a bag-level representation or (2) directly aggregates instance-level scores such as logits, MIL methods may fall into two categories: bag-level and instance-level approaches (see [3]). While many methods from earlier MIL research are instance-level, e.g., [43, 37], later research typically focused on bag-level methods, e.g., [22, 36, 35]. More recently, mainstream deep MIL research has been dominated by attention-based methods [17, 21, 30], which can be viewed as both bag-level and instance-level methods. The attention scores can be viewed as a weighting function that not only aggregates instance-level representations into bag-level ones, but also aggregates instance-level scores. These methods are the focus of our study and are used to derive the benchmark models.
@@
There are also sub-domains of MIL that indirectly relate to our paper. One strand of research considers dynamic representations of bags and models the context of instances [45, 22, 39]. Unlike our study, the focus of this strand is to model the contextual information within a bag, rather than information from outside. Another strand emerging more recently considers the interpretability of weakly supervised learning, proposing evaluation metrics that do [18, 5] or do not [12] require (unobservable) instance labels. We focus on the former metrics in this paper.
@@
Like recent MIL [21, 30] and other [11, 28] research, our methods are also strongly inspired by the transformer architecture [34] and can be grouped with the wider range of research that is based on this seminal model. Indeed, many CV or NLP problems solved by transformers can be represented as MIL problems with sequentially ordered instances, in which the sequence order can be addressed by the addition of positional embedding (see, e.g., [30]).
@@
Finally, there are specific areas of CV and NLP that are related to our work. Examples in CV include weakly-supervised object detection/localization [18, 5] and one-shot object detection [15]. In NLP, one area is “fact extraction and verification” (“FEVER”, see [33]). It is well-suited to the MI verification framework because of the framework’s ability to deal with ambiguity. Indeed, we evaluate our approach on one of the FEVER datasets.
@@
3 MI verification: problem definition and an interpretation of attention scores
@@
The scenario of “verification with noise” is applicable to many tasks in the areas of CV and NLP. To give a few examples, practical use cases of MI verification include:
@@
  • verification of whether a document (bag of paragraphs) has content generated by a particular writer based on a writing sample (paragraph) of this writer;
  • verification of whether a video clip (bag of images) contains targets such as people, endangered species, plants, etc., specified by a reference image;
@@
  • verification of whether a medical image (bag of patches) indicates certain diseases like Covid-19, cancer, etc., based on a known disease image patch. A generalization of this use case relates to the broader subject of “weakly-supervised one-shot object detection” in the CV domain;
@@
  • verification of whether a network log (bag of scripts or time series) records traces of cyber-attacks based on the footprint of a past attack, and so on and so forth.
It is also worth noting that, while standard information retrieval tasks, i.e., searching for content relating to a query, can usually be addressed by classic verification methods, MI verification relates to non-standard, weakly-supervised “cross-domain” retrieval tasks where content is suitably represented as a bag of multiple and ambiguous instances. For example, the first two examples above—retrieving a document based on a paragraph, or retrieving a video based on an image—are relevant use cases.
3.1 Three example MI verification tasks

### v2

@@ Abstract
We explore multiple instance verification, a problem setting in which a query instance is verified against a bag of target instances with heterogeneous, unknown relevancy. We show that naive adaptations of attention-based multiple instance learning (MIL) methods and standard verification methods like Siamese neural networks are unsuitable for this setting: directly combining state-of-the-art (SOTA) MIL methods and Siamese networks is shown to be no better, and sometimes significantly worse, than a simple baseline model. Postulating that this may be caused by the failure of the representation of the target bag to incorporate the query instance, we introduce a new pooling approach named “cross-attention pooling” (CAP). Under the CAP framework, we propose two novel attention functions to address the challenge of distinguishing between highly similar instances in a target bag. Through empirical studies on three different verification tasks, we demonstrate that CAP outperforms adaptations of SOTA MIL methods and the baseline by substantial margins, in terms of both classification accuracy and the ability to detect key instances. The superior ability to identify key instances is attributed to the new attention functions by ablation studies. We share our code at https://github.com/xxweka/MIV.
Keywords: Multiple Instance Learning, Verification, Siamese Neural Networks, Attention, Key Instance Detection
@@ 1 Introduction
In multiple instance (MI) verification, each exemplar for machine learning consists of a pair of two objects—a query instance and a bag of target instances—and a binary class label that indicates whether or not the bag contains an instance pertaining to the same (unobservable) class as the query instance. [^1] For brevity, we will refer to these components as “query”, “target bag”, “exemplar label”, and “query class”, respectively. Figure 1(a) shows an example: a synthetic task based on an extended version of the MNIST data consisting of digits and their writers’ ID (Yadav and Bottou (2019)). In this task, each exemplar contains multiple instances of the same digit written by different writers. The question is whether the target bag contains handwritten digits from the writer who wrote the query—and if there are any, which ones are “key instances”. [^2] We use the terms “key instance” and “witness instance” interchangeably throughout this paper, following standard practice in the literature. Other practical applications are illustrated in Figures 1(b) (signature verification) and 1(c) (fact verification) respectively. Such tasks are of practical importance because they represent settings that require “verification with noise”. While classic verification tasks involve comparing a query instance to a single target instance, real-world targets available for verification are often “noisy” in the sense that they consist of multiple candidates, some irrelevant, to be compared. Importantly, which individual target instances are relevant is unknown, possibly because obtaining labels indicating relevance is too costly. Hence, verification needs to be conducted using a bag of noisy target objects. Note that this generally involves establishing candidate instance relevance to the outcome of the verification process as well—a potentially very useful byproduct.
@@
Conceptually, the MI verification (MIV) problem may be characterized as a combination of two sub-problems: verification (Bromley et al. (1993)) and multiple instance learning (MIL Dietterich et al. (1997)). Verification is frequently formulated as a binary classification task based on a pair of inputs, predicting whether or not the paired inputs are of the same class. The standard solution is to apply Siamese neural networks, as in (Bromley et al. (1993); Chopra et al. (2005); Koch et al. (2015)), but they do not have a built-in mechanism to aggregate or select multiple elements within an input, which makes them unsuitable for MIV. On the other hand, algorithms for MIL, which classify bags of instances, address the ambiguity in how a bag-level class label is related to the unknown labels of the instances in a bag (Zhou (2017); Carbonneau et al. (2018)). However, existing MIL methods do not provide a dedicated mechanism to take into account the information provided by a query instance. In this paper, we argue that new algorithms and models are required for the MIV problem setting, and provide theoretical and empirical evidence showing the benefit of explicitly incorporating both components—the target bag and the query instance—and modeling their relationship during learning and inference. We revisit the distinction between MIV, standard verification tasks, and MIL in Section 3.5.
@@
We also stress that “multiple instance verification” and similar terms used in domains such as biometrics (Liu et al. (2019); Tarek et al. (2022)) refer to problems that differ from the setting considered in our paper. In these problems, all instances of a bag, by definition, have homogeneous unobservable labels, implying that there is no noise or irrelevant elements in the bag for verification. This is in contrast to the verification problem with heterogeneous instances studied in this paper. Crucially, just as in traditional MIL, an exemplar label only provides information about the existence of key instances in a bag but does not identify them. Moreover, since the query instance may belong to a class that has not been observed when the verification model was trained, it enables classification into previously unknown classes. Following the rationale of Koch et al. (2015), this is analogous to an MI version of zero-shot classification.
@@
A naive approach to MIV is to combine existing verification and MIL methods as building blocks, for example, by creating a bag-level representation for the target bag using a state-of-the-art (SOTA) MI pooling method, and then treating this bag-level representation together with the query instance as the two inputs to a Siamese network for verification. Suitable SOTA MI pooling methods include approaches based on gated attention (“Gated-Attention-Based MIL” Ilse et al. (2018)), the transformer decoder (“Pooling by Multi-head Attention” Lee et al. (2019)), and the transformer encoder (e.g., Shao et al. (2021)). Alternatively, non-attention-based methods such as MI-Net (Wang et al. (2018)) and bi-directional LSTM (Bi-LSTM Wang et al. (2020)) can be applied. However, we show that solutions based on this straightforward approach, regardless of which of the above MIL algorithms is applied, yield unsatisfactory classification performance. More specifically, a simplistic “baseline” model is shown to outperform the resulting instantiations—henceforth referred to as “benchmarks”—of the naive approach to MIV.
@@
We postulate that the benchmarks perform unsatisfactorily because they do not incorporate the query instance into the bag-level representation of the target bag: the “query” concept does not exist in the underlying MIL methods. Indeed, we show theoretically, using a probabilistic interpretation of attention scores, that failure to incorporate the query instance induces less effective attention scores. Hence, inspired by the “query-key-value” attention mechanism of the transformer decoder by Vaswani et al. (2017), we propose a new pooling approach called “cross attention pooling” (CAP) that creates the bag-level representation of a target bag through a cross-reference attending to the query. This means that the bag’s representation is dynamic, because it varies when a different query instance is present, even if the bag itself does not change. In contrast, standard MIL methods create a static bag-level representation. CAP also enables the introduction of two novel attention functions, namely “distance-based attention” and “variance-excited multiplicative attention”, respectively, to address the challenge that the instances in a bag may be difficult to distinguish. Both functions accommodate multi-head formulations popular in transformers.
Despite being inspired by “transformer-style” attention, there are significant novelties in CAP’s cross-attention mechanism that cannot be achieved by the standard attention mechanism of transformers. In particular, CAP’s cross-attention facilitates two outputs, by sharing parameters between the two processes producing the outputs alongside attention scores. It also allows for a “co-excitation” mechanism that yields consistent improvements in performance (cf. Section 6.4). Finally, CAP incorporates new attention functions accounting for in-bag, channel-wise variations, not seen in standard transformers’ attention. All these designs make CAP’s cross-attention a bespoke mechanism tailored to MIV, and well-suited for MI pooling. Importantly, simply adapting the commonly used “transformer-like” cross attention—a mechanism intended to construct contextualized embeddings—to the MIV setting (cf. Section 4.2.3) may fail to effectively tackle this problem, even producing unintended results, see Section 6.
@@
In addition to achieving high classification accuracy, the ability to identify key instances in a target bag is also highly desirable for MIV. This ability is sometimes referred to as “explainability” or “interpretability” in the literature on MIL, including (Ilse et al. (2018); Shao et al. (2021); Early et al. (2022)). Following recent related work (Jia et al. (2021); Choe et al. (2023)), we quantitatively evaluate key instance detection for the methods in our study. This is possible for datasets with ground-truth identifiers of key instances.
@@
Our experimental results, obtained from three different tasks and considering both classification and key instance detection, show that CAP is an effective solution to the MI verification problem. In terms of (bag-label) classification performance, it significantly outperforms the benchmarks and the baseline by large margins. Furthermore, using quantitative measures of key instance detection performance, we show that it identifies key instances much more accurately than both the benchmarks and the baseline. Higher classification and key instance detection accuracy are observed with different numbers of training exemplars and with varying bag sizes in the training data. We also present an ablation study showing that the components of the novel attention mechanisms responsible for delivering superior quality of key instance detection—and thus leading to better bag-level representations—are the source of better classification performance.
@@
Our main contributions are as follows. First, we formally state the multiple instance verification problem, and create a new solution, CAP, that outperforms SOTA MIL methods combined with Siamese networks and a simple baseline method on three tasks from the computer vision (CV) and natural language processing (NLP) domains. Furthermore, we propose two new attention functions for CAP: (a) distance-based attention, which differs from the mainstream functional forms of attention because it is neither additive nor multiplicative, and (b) variance-excited multiplicative attention, which is a new form of multiplicative attention. We show that they are better able to identify key instances in MIV, resulting in better bag-level representations. While attention-based methods are not uncommon in MIL-related domains, research studying different types of attention mechanisms and analyzing their efficacy under specific conditions is rare. Within this context, we provide insights that highlight the importance of quantitative evaluation of the ability to detect key instances, which helps identify spurious model performance. We find that MI models may identify key instances incorrectly but still classify the corresponding bag correctly, and vice versa. Our work also contributes to the specific application areas we consider. For example, in the natural language inference (NLI) subtask of FEVER, the existing approach is a two-step pipeline: evidence retrieval and support verification. Our approach combines them into a single step through CAP, and thus enables end-to-end training of the whole pipeline.
@@ 2 Related work
The most closely related existing work to multi-instance verification (MIV) is research on classic verification using Siamese neural networks and the deep MIL literature. The relationships between MIV tasks and problems tackled using learning approaches from these two areas are discussed further in Section 3.5.
@@
The verification problem, and its solution using Siamese networks, was first introduced by the pioneering work of Bromley et al. (1993). Chopra et al. (2005) further explored Siamese networks for learning similarity metrics. Since then, verification problem settings have been generalized in both CV (e.g., Zagoruyko and Komodakis (2015)) and NLP (e.g., Conneau et al. (2017)) domains where Siamese networks can be applied. There are also several areas in machine learning that apply Siamese networks to solve problems similar to verification, including one-shot learning (Koch et al. (2015); Hsieh et al. (2019)), self-supervised contrastive learning (Chen et al. (2020); Zbontar et al. (2021)), embedding pre-training (Conneau et al. (2017); Reimers and Gurevych (2019)), to name a few examples.
@@
We note that there exist approaches other than Siamese networks that may also be applied to MIV in specific application scenarios, such as using concatenation of the query instance and the target bag in an NLP task considered in Sathe and Park (2021). We do not consider this approach here because it needs to be used jointly with other NLP techniques and cannot be trained end-to-end. Within the scope of this paper, we only consider Siamese networks, which provide a more generic and domain-agnostic approach.
@@
Regarding MIL, whose study began with the seminal work of Dietterich et al. (1997) and which can be viewed as a form of “weakly supervised learning” (see Zhou (2017)), there are strong connections to our work. Loosely speaking, depending on whether a method (1) explicitly creates a bag-level representation or (2) directly aggregates instance-level scores (or logits), MIL methods may fall into two categories: bag-level or instance-level approaches (see Carbonneau et al. (2018)). While many methods from earlier MIL research are instance-level (e.g., Zhang and Goldman (2001); Xu and Frank (2004)), later research typically focused on bag-level methods (e.g., Li and Vasconcelos (2015); Wang et al. (2018, 2020)). More recently, mainstream deep MIL research has been dominated by attention-based methods (including Ilse et al. (2018); Lee et al. (2019); Shao et al. (2021)), which can be viewed as both bag-level and instance-level methods. The attention scores can be viewed as a weighting function that not only aggregates instance-level representations into bag-level ones, but aggregates instance-level scores as well. These methods are the focus of our study and are used to derive the benchmark models we use.
@@
There are also sub-domains of MIL that indirectly relate to our paper. One strand of research considers dynamic representations of bags based on modeling the context of instances (Zhou et al. (2009); Li and Vasconcelos (2015); Yan et al. (2018)). Unlike our study, this research focuses on modeling contextual information within a bag, rather than information from outside. Another strand emerging more recently considers the quality of key instance detection in weakly supervised learning, proposing evaluation metrics that do (Jia et al. (2021); Choe et al. (2023)) or do not (Early et al. (2022)) require unobservable instance-level labels. We focus on the former metrics in our study and use them to measure the attention scores’ ability to identify key instances. Our goal is to quantitatively evaluate these scores and draw insights in the context of our theoretical propositions in Section 3.4.
@@
Like recent MIL (Lee et al. (2019); Shao et al. (2021)) and other research (Dosovitskiy et al. (2021); Reimers and Gurevych (2019)), our methods are also strongly inspired by the transformer architecture (Vaswani et al. (2017)) and can be grouped with the wider range of research that is based on this seminal model. Indeed, many CV or NLP problems solved by transformers can be represented as MIL problems with sequentially ordered instances, in which the sequence order can be addressed by the addition of a positional embedding (see, for example, Shao et al. (2021)).
@@
Finally, there are specific areas of CV and NLP that are related to our work. Examples in CV include weakly-supervised object detection or localization (Jia et al. (2021); Choe et al. (2023)) and one-shot object detection (Hsieh et al. (2019)). In NLP, one area is “fact extraction and verification” (FEVER, Thorne et al. (2018)). It is well-suited to the MIV framework because of the framework’s ability to deal with ambiguity. Indeed, we evaluate our approach on one of the FEVER datasets.
@@
3 Multi-instance verification (MIV): problem definition and an interpretation of attention scores
@@
The scenario of “verification with noise” (or “verification with unobservable instance labels”) is applicable to many tasks in the areas of CV and NLP. To give a few examples, practical use cases of MIV may include:
@@
  • verification of whether documents (bags of unlabeled paragraphs) contain content generated by a particular writer based on the query representing a writing sample of this writer;
  • verification of whether bags of unlabeled images contain targets such as people, endangered species, plants, etc., specified by a reference image (query). This use case can be generalized to bags of sequentially ordered images, like video clips; if the sequence order is essential, it can be addressed by the addition of positional embeddings;
@@
  • verification of whether histology whole slide images (bags of unlabeled patches) indicate certain diseases based on a query image of diagnosed (tumor) cells. A generalization of this use case relates to the broader subject of “weakly-supervised one-shot object detection” in the CV domain;
@@
  • verification of whether network logs (bags of unlabeled scripts or time series) record traces of cyber-attacks based on the footprint of a past attack (query), and so on and so forth.
It is worth noting that standard information retrieval tasks, i.e., searching for content relating to a query, can usually be addressed by classic verification methods, without the MIV formulation. However, when the ground truth for standard retrieval tasks is unavailable, that is, it cannot be easily defined whether a unit of searched content “matches” the query, the standard training procedures may be impermissible due to lack of class labels. Under those non-standard circumstances, if a label is available for a bag of (unlabeled) units of content—i.e., a label of whether the bag contains a “match” while the precise match is unknown—MIV may be applicable. For example, the first two examples above—document retrieval based on the writing style of a sample, or searching videos based on a single image of an object—are relevant use cases. Hence, in terms of practical applications, MIV and standard retrieval problems apply to mutually exclusive scenarios, thus complementing each other. The appropriate formulation should be adopted according to available data.
3.1 Three example MIV tasks